The refractive index of air w.r.t. glass is 2/3. The refractive index of diamond w.r.t. air is 12/5. Then the refractive index of glass w.r.t. diamond will be?

Grade:12th pass

1 Answers

shaik omar farooq
5 years ago
Refractive index of air w.r.t to glass is 2/3=mu1 
Refractive index of diamond w.r.t air is 12/5=mu2
Refractive index of glass w.r.t diamond=refractive index of air w.r.tdiamond/refractive index of air w.r.t glass
 
mu2/mu1=mu
mu=5/12/2/3
mu=5/8